### v3

@@ 3.4 Initial action
This $n$ is the subsistence-proof size of potential generations.
Finally, for AK models, the optimal initial action by (19) is $c_{0}^{*}[\infty]=(1-\beta)Ak_{0}$, but this value approaches zero if future generations are given equal consideration as the current generation ($\beta\to 1$). That is, in AK models with an infinite planning horizon, giving ultimate consideration to future generations requires the current generation to ultimately reduce their consumption, i.e., $c_{0}^{*}[\infty]\to{\nu}$. Alternatively, if we relax the infinite horizon assumption, the initial action for AK models with ultimate consideration for the future generations becomes:
@@ 4 Concluding Remarks
The extent to which we account for the entire potential generation upon creating the objective function of the population (i.e., the dynasty) is usually determined by the compounding discounting factor. The potential number of intrinsic years that are accounted for by compound discounting at a rate of $r$ is as high as $\int_{0}^{\infty}e^{-{r}x}\mathrm{d}x=1/r$. With respect to our problem, compound discounting at a rate of $r$ is equivalent to binary discounting (which eliminates consideration for utilities before the planning horizon while fully discounting those that lie beyond it) at a planning horizon $n=1/r$. Binary discounting is flexible in the sense that the planning horizon is variable so that infinity is not precluded as an option. Moreover, it is nondiscriminatory, ensuring that existing generations are treated in ultimate equity. With binary discounting, however, the roundaboutness of production requires a finite horizon dynamic programming approach.
One important feature of our welfare evaluation on the basis of critical-level utilitarianism is that it hinders future generations that earn below-critical-level utility. If the phase boundary $\log(A\beta)$ for the AK setting is negative, the infinite horizon solution trajectory of utility (less the critical level) is forever declining (as shown in the middle-right panel of Figure 3) and eventually surpasses the level of well-being subsistence. Binary discounting avoids such repugnancy by limiting the longevity of the dynasty. On the other hand, if the phase boundary $\log B$ for ZD is negative, a finite horizon must be optimal since an infinite horizon objective function approaches negative infinity. Note also that the bottom utilities that lie at the center of the planning horizon (as shown in the middle-right panel of Figure 4) yield increasingly negative contributions to the population value.
With respect to the distribution of consumption among generations, the rate of compound discounting plays a key role. That is, the lower the rate of discounting is, the more equal the consumption becomes. Moreover, as the planning horizon expands, the distribution of consumption becomes more uneven in an AK setting, whereas vice versa in a ZD setting. Another concern was the optimal initial consumption in light of the well-being of future generations. In any case, the optimal initial consumption decreases in the planning horizon, and ultimately, it depends on the value of $\beta\theta$. If $\beta\theta\to 1$ (that is, with the AK/ZD setting), the optimal initial consumption approaches the level of ultimate sacrifice under an infinite planning horizon, Conversely, any finite consumption of the current generation determines the (finite) planning horizon, i.e., the size of future generations.
@@
where $\rho>0$ denotes the rate of interest. Hence, if $\theta=1$ (AK setting), $A=B=\text{MPK}=\rho+\delta$, and if $\delta=1$ (complete depreciation), it must be the case that $A=B=1+\rho>1$.
From the study in Section 3.1, we note that the condition $A\beta<1$ implies the nonincreasing property of undiscounted contributions (or utilities less the critical level) with respect to generations. The knife-edge case where $A\beta=1$ implies that the undiscounted contributions are constant over generations. Figure 7 (left) depicts the set of possible values for $(\theta,A)$ that satisfies $A\beta<1$. Additionally, from the study in Section 3.2, we note that the condition $B<1$ implies the existence of a finite optimal planning horizon. Otherwise, if $B>1$, the optimal planning horizon will be infinite. The knife-edge case where $B=1$ implies that the optimal planning horizon is indeterminate. Figure 7 (right) depicts the set of possible values for $(\theta,A)$ that satisfies $B<1$.
@@
We show this by induction. First, the value function at $t={n}$ is evaluated via (4):
@@
Since $k_{{n}+1}=0$ by (2.1), $c_{n}={A}(k_{n})^{\theta}$ must be true. Additionally, $\mathcal{V}_{{n}+1}=0$ must be true for efficiency. Thus, the final maximization is bounded, i.e.,
@@
Because $S_{n-n}=S_{0}=1$ and since ${A}$ is a constant, the lemma holds true for $t={n}$.
Suppose that the lemma holds true for $t+1$. Then,
@@
We show this by induction. By plugging (21) into (2.1), we obtain:
@@
We know that the lemma is true for $t=1$.
Suppose that the lemma is true for $t$. We then know by (22) that:

### v4

@@ 3.4 Initial action
This $N$ is the subsistence-proof size of potential generations. Finally, for AK models, the optimal initial action by (22) is $c_{0}^{*}[\infty]=(1-\beta)Ak_{0}$, but this value approaches zero if future generations are given equal consideration as the current generation ($\beta\to 1$). That is, in AK models with an infinite planning horizon, giving ultimate consideration to future generations requires the current generation to ultimately reduce their consumption, i.e., $c_{0}^{*}[\infty]\to{\nu}$. Alternatively, if we relax the infinite horizon assumption, the initial action for AK models with ultimate consideration for the future generations becomes:
@@ 4 Concluding Remarks
This paper has investigated the optimal longevity of a dynasty under a Critical-Level Utilitarian framework. By treating the planning horizon as an endogenous choice variable within a dynamic programming setting, we derived closed-form analytical solutions that link economic fundamentals to the ethical duration of a society. Our analysis highlights two distinct regimes determined by productivity and time preference. In the AK setting with discounting, the sign of $\log(A\beta)$ serves as the conservative boundary determining whether the optimal horizon is finite or infinite. Similarly, in the Zero Discounting (ZD) setting, which prioritizes intergenerational equity, the sign of the cost-based productivity $\log(B)$ dictates the optimal longevity. These results demonstrate that under low productivity or strict ethical constraints, a finite horizon is not an anomaly but a structurally optimal response to avoid negative social welfare.
Our findings also reveal a striking contrast in intergenerational inequality between the two settings. In the AK setting, a longer planning horizon is associated with greater inequality across generations, although a higher discount factor $\beta$ mitigates this effect. Conversely, in the ZD setting, extending the horizon tends to reduce inequality, as the burden of capital accumulation is shared more evenly. However, a higher output elasticity of capital $\theta$ in the ZD model exacerbates long-term inequality. These observations suggest that the pursuit of longevity imposes different distributional costs depending on the underlying technology and ethical discount rates.
A key theoretical insight emerges from the behavior of the initial consumption $c_{0}$ as we approach the limit of perfect intergenerational impartiality ($\beta\to 1$). In both models, a higher $\beta$ generally requires the current generation to reduce consumption to support a longer dynasty. In the limiting case of the infinite-horizon AK model with $\beta\to 1$ (and $\theta=1$), the optimal initial consumption approaches zero (or the subsistence level). This implies that if the future holds infinite potential, the current generation is ethically compelled to make the ultimate sacrifice for the sake of an eternal dynasty. In contrast, in a stagnant economy where such sacrifice yields diminishing returns, our model prescribes termination. Thus, our framework illuminates an ethical asymmetry: it demands immense endurance when the future is promising, but offers the “mercy” of termination when the future promises only misery.
Methodologically, our approach ensures time consistency through the recursive structure of dynamic programming. Unlike ad hoc termination rules, the finite horizon derived here is the result of a consistent optimization process where the continuation value falls below the critical threshold. We acknowledge, however, that our model relies on a logarithmic utility function and a fixed population size per generation to maintain analytical tractability. Introducing more general preferences (such as CRRA utility), endogenous population size, or environmental constraints would likely preclude closed-form solutions and require numerical approaches. Exploring these complex dynamics remains a promising avenue for future research.
Finally, our results offer a normative perspective on the demographic trends observed in modern mature economies. The choice to limit the number of offspring—or in our macroeconomic interpretation, to limit the longevity of the dynasty—is often viewed as a failure of vitality. However, from the perspective of Critical-Level Utilitarianism, this may be interpreted as an altruistic termination. If the current generation anticipates that economic or environmental constraints will condemn future generations to lives barely worth living, choosing a finite horizon is a rational and ethical decision to prevent the creation of suffering.
@@
where $\rho>0$ denotes the rate of interest. Hence, if $\theta=1$ (AK setting), $A=B=\text{MPK}=\rho+\delta$, and if $\delta=1$ (complete depreciation), it must be the case that $A=B=1+\rho>1$. From the study in Section 3.1, we note that the condition $A\beta<1$ implies the nonincreasing property of undiscounted contributions (or utilities less the critical level) with respect to generations. The knife-edge case where $A\beta=1$ implies that the undiscounted contributions are constant over generations. Figure 7 (left) depicts the set of possible values for $(\theta,A)$ that satisfies $A\beta<1$. Additionally, from the study in Section 3.2, we note that the condition $B<1$ implies the existence of a finite optimal planning horizon. Otherwise, if $B>1$, the optimal planning horizon will be infinite. The knife-edge case where $B=1$ implies that the optimal planning horizon is indeterminate. Figure 7 (right) depicts the set of possible values for $(\theta,A)$ that satisfies $B<1$.
@@
We show this by induction. First, the value function at t=Nt={N} is evaluated via (7):
@@
Since $k_{{N}+1}=0$ by (6b), $c_{N}={A}(k_{N})^{\theta}$ must be true. Additionally, $\mathcal{V}_{{N}+1}=0$ must be true for efficiency. Thus, the final maximization is bounded, i.e.,
@@
Because $S_{N-N}=S_{0}=1$ and since A{A} is a constant, the lemma holds true for t=Nt={N}. Suppose that the lemma holds true for $t+1$. Then,
@@
We show this by induction. By plugging (24) into (6b), we obtain:
@@
We know that the lemma is true for $t=1$. Suppose that the lemma is true for $t$. We then know by (25) that: